OpenStudy (anonymous):

If prices are reduced 25% and sales increase 20%, what is the net effect on gross receipts? What is the net effect? and how do you get it?

OpenStudy (valpey):

Reduced by 25% is the same a multiplied by 0.75. Increased by 20% is the same as multiplied by 1.20. Multiply the two multipliers together to get the net effect.

OpenStudy (anonymous):

how is increased by 20% is same as multiplied by 1.2? and one % is on the sale and the other is on the price itself, how can I work with a problem like that?

OpenStudy (valpey):

If I sell 100 things a day and sales go up 20% I sell 120 things a day. That is the same as multiplying sales by 1.20. Gross Receipts = Sales Price * Number of units sold. If I change Sales Price and Number of units sold each by simple multipliers, the effect on Gross Receipts will be the product of the multipliers.

OpenStudy (valpey):

Let GR = Gross Receipts; S = Sales; and P = Price. Then: GR = S * P Now lets change price to 0.75*P and lets change S to 1.20*S. Now, let NGR = New Gross Receipts and: NGR = 0.75*P * S to 1.20*S

OpenStudy (valpey):

Oops, NGR = 0.75*P * 1.20*S

OpenStudy (anonymous):

perfect!!! now what?

OpenStudy (anonymous):

how do we get the effect?

OpenStudy (valpey):

Multiply the Multipliers together. This will give you the net multiplier.

OpenStudy (anonymous):

soit's 90%?

OpenStudy (valpey):

New gross receipts will be 90% of old gross receipts, yes. You might also say "Gross Receipts will go down by 10%"

OpenStudy (anonymous):

ok thanks

OpenStudy (valpey):

Cheers